What function best describes the sequence ( 7,-1,-9,-17,...)? I think it is tn= -8n+1 but I'm not sure . Can someone please explain this to me? Thanks.
Answer by mayank(15)

Ans. As the sequence is (7,-1,-9,-17...) the first term clearly is 7.
Let it be t1.thus t1=7
if we let second term as t2 then t2=-1 and if we let third term as t3 then t3=-9.
Let nth term be tn.
Now considering the sequence as arithmetic sequence, we know that for any arithmetic sequence the nth term is given by tn=t1+(n-1)d, where d is constant difference between consecutive terms.
Thus here tn=7+(n-1)-8[As difference d here is -8,t2-t1=-1-7=-8]
thus tn=7-8(n-1)
thus tn=7-8n+8
thus tn=15-8n.
Thus ans is tn=15-8n. Thus this function best describes the above sequence.
[Note: here we have considered the sequence to be arithmetic one]